NEED HELP ASAP!!!
rectangle with length 6 meters more than half it's width. The perimeter is 24 meters. Find the length and width

Answer by mananth(16949)

width = w
length = w/2 +6
...
Perimeter = 2(l+w)
24= 2(w+(w/2 +6))
24=2(w+(w+12)/2)
24=2(2w+w+12)/2
24=3w+12
3w=12
/3
w=4
L= 8